मराठी
सी.आई.एस.सी.ई.आयसीएसई ICSE Class 7

Five Students A, B, C, D, and E Are Competing in a Long-distance Race. Each Student’S Probability of Winning the Race is Given Below: a → 20%, B → 22%, C → 7%, D → 15% and E → 36%

Advertisements
Advertisements

प्रश्न

Five Students A, B, C, D, and E are competing in a long-distance race. Each student’s probability of winning the race is given below:
A → 20%, B → 22%, C → 7%, D → 15% and E → 36%
(i) Who is most likely to win the race?
(ii) Who is least likely to win the race?
(iii) Find the sum of probabilities given.
(iv) Find the probability that either A or D will win the race.
(v) Let S be the event that B will win the race.
(a) Find P(S)
(b) State, in words, the complementary event S’.
(c) Find P(S’)

बेरीज
Advertisements

उत्तर

Given probabilities of five students A, B, C, D, and E such as
P(A) = 20%, P(B) = 22%, P(C) = 7%, P(D) = 15%, and P(E) = 36%

(i) The mostly chance of winning the race is of Student E. ..........[∵ P(E) = 36% maximun]

(ii) The least chances of winning the race is of Student C. ..........[∵ P(C) = 7% minimum]

(iii) The sum  of the probabilities

= P(A) + P(B) + P(C) + P(D) + P(E) 

= 20% + 22% + 7% + 15% + 36%

= 100%

(iv) Favourable outcomes that either A or D will win = 20% + 15% = 35%

P(either A or D will win) =`35/100=7/20`

(v) (a) Favourable outcomes that B will win = 22%

P(S) =`22/100=11/50`

(b) S' = B will not win the race.

(c) P(S') = 1 − P(S)

= `1-11/50=(50-11)/50=39/50`

shaalaa.com
Concept for Feel of Probability Using Data Through Experiments.
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 25: Probability - Exercise 22 (B)

APPEARS IN

संबंधित प्रश्‍न

In a cricket match, a batsman hits a boundary 12 times out of 80 balls he plays, further, if he plays one ball more, what will be the probability that:
(i) he hits a boundary
(ii) he does not hit a boundary


There are 8 marbles in a bag with numbers from 1 to 8 marked on each of them. What is the probability of drawing a marble with a number
(i) 3
(ii) 7


Two coins are tossed simultaneously 100 times and the outcomes are as given below:

Outcomes Two heads
(H, H)
Exactly one head
(H T or T H)
No head
(T T)
No. of times 21 55 24

If the same pair of coins is tossed again at random, find the probability of getting:
(i) two heads
(ii) exactly one head
(iii) no head.


A bag contains 4 white and 6 black balls,- all of the same shape and same size. A ball is drawn from the bag without looking into the bag. Find the probability that the ball drawn is:
(i) a black ball
(ii) a white ball
(iii) not a black ball


In a single throw of a dice, find the probability of getting a number 4


In a single throw of a dice, find the probability of getting a number 6


Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 50


Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is 40


A carton contains eight brown and four white eggs. Find the probability that an egg selected at random is:
(i) brown
(ii) white


A box contains 3 yellow, 4 green, and 8 blue tickets. A ticket is chosen at random. Find the probability that the ticket is:
(i) yellow
(ii) green
(iii) blue
(iv) red
(v) not yellow


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×